OpenStudy (anonymous):

Which rate is the best deal for a package of peanuts? A. $1.40 for 50 g B. $1.02 for 35 g C. $2.25 for 80 g D. $2.05 for 75 g

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

this is my question idk

OpenStudy (anonymous):

which one do you tihnk is your answer love

OpenStudy (anonymous):

a or b?

OpenStudy (dtan5457):

Wouldn't you just take a common multiple of all your grams, multiply them accordingly, and see which is the cheapest.

OpenStudy (anonymous):

@dtan5457 so would A be the answer

OpenStudy (anonymous):

@dtan5457 Or maybe D?

jimthompson5910 (jim_thompson5910):

divide each price by the number of grams to get the unit price then compare each unit price (smallest unit price wins)

OpenStudy (anonymous):

Im thinking D

jimthompson5910 (jim_thompson5910):

1.4/50 = 0.028 1.02/35 = 0.02914285714286 2.25/80 = 0.028125 2.05/75 = 0.02733333333333 D is the best deal. I agree

OpenStudy (anonymous):

thank you